40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11919     Entry time: Thu Jan 7 16:52:32 2016     In reply to: 11917     Reply to this: 11920
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX X Freq Noise measured 

Here is some of the promised data. As mentioned, changing diode current and crystal temperature didn't have much effect on the frequency noise spectrum; but the spectrum itself does seem too high for our needs. 

At each temperature, we started measuring the spectrum at 1.8A, and stepped the current up, hoping to reach 2.0 A.

At 47.5 C, we were able to scan the current from 1.8 to 2.0 A without much problem. At 49.0C, the laser mode would hop away above 1.95A. At 50.4C it would hop away above 1.85A. The spectra were not seen to change when physically disconnecting the PZT actuation BNC from the rear of the laser. 

The flattening out at the upper end is likely due to the SR560 output noise. I foolishly neglected to record the output spectrum of it, but with the marconi external modulation set to 3.2MHz/V, the few Hz/rtHz above 20k translates to a signal on the order of uV/rtHz, which seems reasonable. 

Data and code attached. 

Attachment 1: AUXfreqnoise.pdf  65 kB  | Hide | Hide all
AUXfreqnoise.pdf
Attachment 2: auxXmeasurements.zip  87 kB
ELOG V3.1.3-